Write the complex number e^i5pi/6 in the form a+bi. Help would truly be appreciated to understand this. I looked around for ways to solve and still cant get it.
Found 3 solutions by Fombitz, stanbon, Alan3354:
Answer by Fombitz(32388)
Answer by stanbon(75887)

Write the complex number e^i5pi/6 in the form a+bi. Help would truly be appreciated to understand this. I looked around for ways to solve and still cant get it.
---------------
Use Google to search for e^i(theta)
---------------------------------------------
e^i(5/6)pi = cos(5/6pi) + isin(5/6pi)
----
= -sqrt(3)/2 + i(1/2)
---
Cheers,
Stan H.
------------------

Answer by Alan3354(69443)

You got 2 answers from 2 tutors and the answers match.
----
It's more interesting to get 2 different answers, and you actually have to use your brain to figure out which is the correct answer.
---------
For now, you got others to do it for you.
Might work out in life, might not.